I’m a certain Simple harmonic motion the projection of a point rotating at 500rev/min has an acceleration of 35m/s2 when the velocity of the point is 5m/s at the initial instant t=0

calculate the frequency of the oscillation in Ho


1
Expert's answer
2021-01-16T17:20:55-0500
"f=\frac{\omega}{2\pi}\\f=\frac{500}{2\pi}\frac{2\pi}{60}=8.33\ Hz"